OpenStudy (anonymous):

order from least to greatest. 1.5, 1/5, 50%, and 0.05

OpenStudy (pixiedust1):

Okay! We see that these are all in different forms. To compare, let's make them all percents! To do that, we multiply everything that is a decimal by 100. 1.5*100=150% 0.05*100=5% Now, 50% is already a percent. We can leave that alone. In order to get a fraction into a percent, we divide 1 by 5, to get 0.2 0.2*100=20% Now, this is easy to order! We have 5%, 20%, 50%, and 150% Therefore, it's 0.05, 1/5, 50%, and 1.5

OpenStudy (pixiedust1):

I hope that helps! :D

OpenStudy (pixiedust1):

@taylee1 Open Study values the learning process. please do not give away answers.

OpenStudy (pixiedust1):

Code of conduct " Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!"

OpenStudy (anonymous):

thank you so much guys you guys are such help

OpenStudy (pixiedust1):

You are very welcome

OpenStudy (anonymous):

what are you sorry about im so confused

OpenStudy (pixiedust1):

I just wanted to make sure that you still got my explanation, even though taylee gave away the answer :)

OpenStudy (anonymous):

well i still got the explantion so yeah but its fine

OpenStudy (anonymous):

but yeah

OpenStudy (pixiedust1):

okay, all's well that ends well :D

OpenStudy (anonymous):

@PixieDust1 oh shut up! Just leave it alone your freaking idiot!